A car travels at 20 m s-1 for 1 h, and then accelerates at 3 m s-2 for 60 s. What is the distance traveled?


1
Expert's answer
2021-01-20T17:18:41-0500

Let's first find the distance traveled by the in 1 hour:


"d_1=v_1t_1=20\ \dfrac{m}{s}\cdot3600\ s=72000\ m."

Then, we can find the distance traveled by the car in 60 seconds when it accelerates:


"d_2=v_0t+\dfrac{1}{2}at^2,""d_2=20\ \dfrac{m}{s}\cdot60\ s+\dfrac{1}{2}\cdot3\ \dfrac{m}{s^2}\cdot(60\ s)^2=6600\ m."

Finally, we can find the total distance traveled by the car:


"d_{tot}=d_1+d_2=72000\ m+6600\ m=78600\ m=78.6\ km."

Answer:

"d_{tot}=78600\ m=78.6\ km."